# The frame-dragging vector potential on galaxy scales from DM-only Newtonian $N$-body simulations

### v2

@@ 5.2 The gravito-magnetic potential
Again here, the vector power spectrum shows even more clearly the power suppression due to the missing perturbation modes larger than the simulation box, as pointed out in the previous sections for the source fields. Although thomasFullyNonlinearPostFriedmann2015 concluded that there is no evidence for a systematic dependence of the vector potential spectrum on box size for boxes smaller than $200\;h^{-1}$ Mpc (with the smallest considered being of $80\;h^{-1}$ Mpc), a clear box size effect seems already visible in all their plots (thomasFullyNonlinearPostFriedmann2015, , see in particular Figure RB6 in Supplementary Material).
@@
For TNG300 and TNG100, this ratio turns out to be around $1\text{--}3\times 10^{-5}$, in agreement with previous results for $z=0$ (bruniComputingGeneralrelativisticEffects2014a; thomasGravityNonlinearScales2015; adamekGeneralRelativityCosmic2016; barrera-hinojosaVectorModesLCDM2021), as well as with 2nd-order PT (luCosmologicalBackgroundVector2009). However, for the smallest simulation, TNG50, the ratio drops down of an order of magnitude due to the box size effect.
The integral (17) couples different modes of the matter density and velocity potential. This means that at a single scale the vector power spectrum receives contributions from different wave numbers. Therefore, in order to compare the vector power spectra obtained from the simulations with the second-order perturbation theory in a consistent way, when performing the integration, one would, in principle, need to set both large-scale and small-scale cut-offs to account for the limited size of the simulation box and the finite sampling (the latter comprising both the limited mass and spatial resolutions of the simulation), respectively. parkbaghyeonImpactBaryonicPhysics2018 developed a procedure to correct the momentum power spectrum for the missing power in the context of the Sunyaev-Zel’dovich effect prediction, which, to the best of our knowledge, was the first attempt of this kind in the literature. In the present work, in order to retain the perturbation theory prediction as a reference, we apply this procedure to the power spectrum of the vector potential obtained from the simulations. The current paper is also the first one applying the parkbaghyeonImpactBaryonicPhysics2018 method to correct the missing power in the gravitomagnetic potential power spectrum.
@@
The resulting power spectrum containing the missing power is then added to the measured power spectrum of the vector potential. In the linear regime, this corrected power spectrum should match the reference second-order perturbation theory prediction $\Delta_{\mathbf{B}}^{\mathrm{PT}}$. Therefore, we checked that the corrected power spectrum actually matches the full power spectrum at $z=20$ (the highest redshift for which the simulation snapshot is available), where the second-order perturbation theory is still expected to hold. Only for the $75\;h^{-1}$ Mpc simulation box, this correction overestimates the reference perturbation theory prediction by 55%, meaning that more power is added than necessary. The same result was found by parkbaghyeonImpactBaryonicPhysics2018 for the Illustris simulation of $75\;h^{-1}$ Mpc, where they calibrated the upper integration limit with a fudge factor $f_{c}$ to compensate for this overshooting, so that $k^{\prime}<f_{c}k_{\mathrm{box}}$. Performing the same calibration, it is found that $f_{c}=1$ for the $205$ and $35$ $h^{-1}$ Mpc boxes, indicating no overestimation, whereas a value of $f=0.9$ is obtained for TNG100-2-Dark.
@@
To conclude, the ratio of the vector and scalar power spectra remains of $\mathcal{O}(10^{-5})$ for the three simulations at $z=0$, confirming both the results found by (bruniComputingGeneralrelativisticEffects2014a) and thomasFullyNonlinearPostFriedmann2015 from Newtonian $N$-body simulations, and those of (barrera-hinojosaGRAMSESNewRoute2020) and (adamekGeneralRelativityCosmic2016) from GR simulations. This means that, up to the scales investigated in this study, i.e. $k\lessapprox 92\;h\,$Mpc-1, which at the present time correspond to a physical scale of $\approx 100$ kpc, the fully non-linear gravito-magnetic potential is around $10^{-2}\text{--}10^{-3}$ times the scalar potential, namely that the frame-dragging is potentially an order $1\text{--}0.1\%$ effect also on the scale of galaxies.
@@ 5.3 Representation of the fields
The density field also shows a clear correlation with the velocity vorticity. The latter is peaked locally in the collapsing regions, where the orbit crossing of the particles occurs more frequently. However, as mentioned in Section 4.1, the velocity divergence and vorticity estimated by DTFE are not completely reliable near caustics (hahnPropertiesCosmicVelocity2015), where velocity gradients are found to diverge. Therefore, such maps only provide qualitative information and an accurate picture on large scales.
@@
A striking feature of the visualization is the emergence of vortical structures in the vector potential, particularly around deeper scalar potential wells. These organised vortical patterns in the vector potential first confirm that the reconstruction of the gravito-magnetic potential, outlined in Section 4.2.1, produced consistent results, as they underline the solenoidal behaviour of the potential. Secondly, they suggest that, in line with the main subject of this paper, while traditional treatments of dark matter assume a predominantly irrotational velocity field, also the rotational dark matter flows may influence the angular momentum transport and thus the evolving cosmic structures in a non-trivial way. In the words of adamekGeneralRelativityCosmic2016, this indicates how spacetime is dragged around by vortical matter flows.
@@ 5.4 Time evolution
Compared to the right panel of Figure 5 in barrera-hinojosaVectorModesLCDM2021, where the power spectrum of the vector potential is found to decrease over time, from $z=1.5$ to $z=0$, the reader should keep in mind that there is a factor $a^{2}$ difference between the adopted conventions for the power spectrum of $\mathbf{B}$. Moreover, in barrera-hinojosaVectorModesLCDM2021, the ratio is found to increase towards the smallest scales, in contrast with our results. This is most likely due to the Adaptive Mesh Refinement technique implemented in GRAMSES, which enables a much higher spatial resolution than DTFE. It is still unclear, however, whether this increase manifests as a result of spurious power introduced by the AMR or rather as a pure GR indication.
Since the non-linear prediction of the vector potential power spectrum $\Delta_{\mathbf{B}}^{\mathrm{n.l.}}$ gives overall a good estimate for the fully non-linear vector potential, and its evolution, relevant information can already be retrieved from this quantity. In order to highlight the redshift dependence of the power spectrum independently of its overall amplitude at a given redshift due to the linear growth of structure, we normalise the power spectrum $\Delta_{\mathbf{B}}^{\mathrm{n.l.}}$ by the analytic growth factor of $\Delta_{\mathbf{B}}^{\mathrm{PT}}$, which is (see Equation B.4 of adamekNbodyMethodsRelativistic2014)
@@
is the growth factor for the scalar potential and $g_{\infty}$ is chosen so that $g(0)=1$ (see Equation A.4 of luCosmologicalBackgroundVector2009). The left panel of Figure 18 shows $\Delta_{\mathbf{B}}^{\mathrm{n.l.}}/f_{\mathbf{B}}^{\mathrm{PT}}$ for different redshifts.
@@
Here, we can clearly see how the non-linear growth of vector modes is highly scale-dependent and how the power in transferred over time. As expected, the non-linearity in the power spectrum of the gravito-magnetic vector potential is observed to increase with redshift from $z=20$ to $z=0$. From the right panel, one can quantify the increase of the amplitude of the non-linear approximation with respect to perturbation theory. In particular, at $z=0$, the non-linear amplitude is $\approx 175$ times larger than PT, for scales close to k=100hk=100\;h Mpc-1, while being $\approx 75$ times larger for k=10hk=10\;h Mpc-1.
@@
For the smallest scales, in the range $10\text{--}100\;h$ Mpc-1, the pure non-linear evolution is intensively boosted from $z=1\text{--}2$ onwards. However, as shown in the right panel and already stated at the beginning of this section, the ratio between the non-linear approximation for the power spectrum of the vector potential and the non-linear analogue for the scalar potential remains of the order of $10^{-5}$. Contrary to thomasFullyNonlinearPostFriedmann2015, where this ratio is not monotonic over the redshift range (see Fig. 9 of the reference), here instead we find a monotonic increase over time. In particular, for the smallest scales, this ratio starts growing rapidly since $z=1\text{--}2$. Therefore, we can conclude that there is no indication pushing towards the direction of a relevant gravito-magnetic effect arising during a particular phase in the evolution of structures, other than the rapidly accelerated yet continuous growth of non-linearity at low redshift.
@@ 6 Conclusions
On non-linear scales, cosmic structure formation is typically studied through Newtonian $N$-body simulations, while relativistic effects are only considered at large scales, of the order of the Hubble horizon, using perturbative methods, e.g. to study primordial non-gaussianity signatures in galaxy surveys. The post-Friedmann approach provides a unified framework for small and large scales that generalises the weak-field approximation to cosmology. In this framework, at the leading order in the $c^{-n}$ expansion, namely in the Newtonian limit of small peculiar velocities, a frame-dragging effect is expected to arise in terms of a gravito-magnetic vector potential. This potential is sourced by the transverse part of the momentum field and is thus expected to be non-zero if measured from Newtonian $N$-body simulations, as confirmed by previous works on cosmological scales (bruniComputingGeneralrelativisticEffects2014a; thomasFullyNonlinearPostFriedmann2015).
The IllustrisTNG project provides a suite of simulations that offer a unique opportunity to investigate this gravito-magnetic effect not only on cosmological scales but also in the highly non-linear regime, down to galactic scales, where structures have fully collapsed, have gone through the virialisation process and the vorticity that has been generated in the multi-stream regime becomes significant (pueblasGenerationVorticityVelocity2009). Our goal here has been the investigation of gravito-magnetic effects on these previosly unexplored smaller non-linear scales, using the post-Friedman framework on simulations of the IllustrisTNG project.
To this aim, three dark-matter-only simulations have been considered, i.e. TNG300-2-Dark, TNG100-2-Dark, and TNG50-2-Dark, which sample different non-linear regimes in the formation of structures. The source fields of the gravito-magnetic potential have been extracted via the Delaunay Tessellation Field Estimator, that provides a detailed reconstruction of the density and velocity fields, from which power spectra were computed.
@@
In our results the power spectra of the velocity gradients show the onset of non-linearity and the generation of vorticity. At around $k\approx 1.5\;h$ Mpc-1 for the $205$ and $75$ $h^{-1}$ Mpc simulations, and at around $k\approx 3\;h$ Mpc-1 for the $35$ $h^{-1}$ Mpc simulation, the velocity divergence shows a dip while the vorticity exhibits a peak and start to dominate at smaller scales as the system transitions into the non-linear regime, in agreement with previous studies (pueblasGenerationVorticityVelocity2009; jelic-cizmekGenerationVorticityCosmological2018; barrera-hinojosaVectorModesLCDM2021). This reveals that rotational (i.e. divergence-free) modes emerge naturally in cosmological simulations as a consequence of orbit crossing in collapsing regions, a process that becomes increasingly important as structures evolve beyond the single-stream approximation due to the collision-less nature of dark matter.
@@
The simulation results are compared with both second-order perturbation theory and a non-linear approximation. This work shows for the first time that the non-linear prediction provides actually a good estimate for the non-linear power spectrum of the vector potential. A correction for the missing power in the gravito-magnetic vector potential power spectrum, inspired by a technique used in the Sunyaev-Zel’dovich effect literature (parkbaghyeonImpactBaryonicPhysics2018) and computed using the non-linear matter power spectrum from HaloFit, confirms that at large scales, the corrected spectra match perturbation theory predictions, while at small scales, the non-linear approximation remains a valid estimate. This also confirms that the correction employed is actually accurate in reproducing consistent results across simulations of different size.
The ratio of the vector to scalar potential, corrected for the missing power, is found to be $\approx 2\text{--}4\times 10^{-5}$ across all simulations, consistent with previous studies (bruniComputingGeneralrelativisticEffects2014a; thomasFullyNonlinearPostFriedmann2015; adamekGeneralRelativityCosmic2016; barrera-hinojosaVectorModesLCDM2021) and second-order perturbation theory (luCosmologicalBackgroundVector2009). Interestingly, the non-linear prediction for the vector power spectrum is found to reach an approximately constant value of $\approx 5\times 10^{-5}$ at very small scales, which is only slightly larger than expected in 2nd-order PT, i.e. $\approx 3\times 10^{-5}$.
Furthermore, by extracting a global solution for the gravito-magnetic potential in the Fourier space, a visual representation of its magnitude is presented for the first time for Newtonian $N$-body simulations. The gravito-magnetic potential, while weaker, is generally correlated with the scalar potential. It also shows clear non-local structures and organised vortical patterns that indicate a coupling between rotational flows and the dragging of spacetime.
@@
WB and MC are indebted to the Italian Space Agency (ASI) for their continuing support through contract 2018-24-HH.0 and its addendum 201824-HH.1-2022 to the National Institute for Astrophysics (INAF). CB-H is supported by the Chilean National Agency of Research and Development (ANID) through grant FONDECYT/Postdoctorado No. 3230512.

### v3

@@ 5.2 The gravito-magnetic potential
Again here, the vector power spectrum shows even more clearly the power suppression due to the missing perturbation modes larger than the simulation box, as pointed out in the previous sections for the source fields. Although Thomas et al. (2015b) concluded that there is no evidence for a systematic dependence of the vector potential spectrum on box size for boxes smaller than $200\;h^{-1}$ Mpc (with the smallest considered being of $80\;h^{-1}$ Mpc), a clear box size effect seems already visible in all their plots (Thomas et al., 2015b, , see in particular Figure RB6 in Supplementary Material).
@@
For TNG300 and TNG100, this ratio turns out to be around $1\text{--}3\times 10^{-5}$, in agreement with previous results for $z=0$ (Bruni et al., 2014; Thomas et al., 2015a; Adamek et al., 2016; Barrera-Hinojosa et al., 2021), as well as with 2nd-order PT (Lu et al., 2009). However, for the smallest simulation, TNG50, the ratio drops down of an order of magnitude due to the box size effect.
The integral (17) couples different modes of the matter density and velocity potential. This means that at a single scale the vector power spectrum receives contributions from different wave numbers. Therefore, in order to compare the vector power spectra obtained from the simulations with the second-order perturbation theory in a consistent way, when performing the integration, one would, in principle, need to set both large-scale and small-scale cut-offs to account for the limited size of the simulation box and the finite sampling (the latter comprising both the limited mass and spatial resolutions of the simulation), respectively. Park et al. (2018) developed a procedure to correct the momentum power spectrum for the missing power in the context of the kSZ effect prediction, which, to the best of our knowledge, was the first attempt of this kind in the literature. In the present work, in order to retain the perturbation theory prediction as a reference, we apply this procedure to the power spectrum of the vector potential obtained from the simulations. The current paper is also the first one applying the Park et al. (2018) method to correct the missing power in the gravito-magnetic potential power spectrum.
@@
The resulting power spectrum containing the missing power is then added to the measured power spectrum of the vector potential. In the linear regime, this corrected power spectrum should match the reference second-order perturbation theory prediction $\Delta_{\mathbf{B}}^{\mathrm{PT}}$. Therefore, we checked that the corrected power spectrum actually matches the full power spectrum at $z=20$ (the highest redshift for which the simulation snapshot is available), where the second-order perturbation theory is still expected to hold. Only for the $75\;h^{-1}$ Mpc simulation box, this correction overestimates the reference perturbation theory prediction by 55%, meaning that more power is added than necessary. The same result was found by Park et al. (2018) for the Illustris simulation of $75\;h^{-1}$ Mpc, where they calibrated the upper integration limit with a fudge factor $f_{c}$ to compensate for this overshooting, so that $k^{\prime}<f_{c}k_{\mathrm{box}}$. Performing the same calibration, it is found that $f_{c}=1$ for the $205$ and $35$ $h^{-1}$ Mpc boxes, indicating no overestimation, whereas a value of $f=0.9$ is obtained for TNG100-2-Dark.
@@
To conclude, the ratio of the vector and scalar power spectra remains of $\mathcal{O}(10^{-5})$ for the three simulations at $z=0$, confirming both the results found by (Bruni et al., 2014) and Thomas et al. (2015b) from Newtonian $N$-body simulations, and those of (Barrera-Hinojosa and Li, 2020) and (Adamek et al., 2016) from GR simulations. This means that, up to the scales investigated in this study, i.e. $k\lessapprox 92\;h\,$Mpc-1, which at the present time correspond to a physical scale of $\approx 100$ kpc, the fully non-linear gravito-magnetic potential is around $10^{-2}\text{--}10^{-3}$ times the scalar potential, namely that the frame-dragging is potentially an order $1\text{--}0.1\%$ effect also on the scale of galaxies.
@@ 5.3 Representation of the fields
The density field also shows a clear correlation with the velocity vorticity. The latter is peaked locally in the collapsing regions, where the orbit crossing of the particles occurs more frequently. However, as mentioned in Section 4.1, the velocity divergence and vorticity estimated by DTFE are not completely reliable near caustics (Hahn et al., 2015), where velocity gradients are found to diverge. Therefore, such maps only provide qualitative information and an accurate picture on large scales.
@@
A striking feature of the visualization is the emergence of vortical structures in the vector potential, particularly around deeper scalar potential wells. These organised vortical patterns in the vector potential first confirm that the reconstruction of the gravito-magnetic potential, outlined in Section 4.2.1, produced consistent results, as they underline the solenoidal behaviour of the potential. Secondly, they suggest that, in line with the main subject of this paper, while traditional treatments of dark matter assume a predominantly irrotational velocity field, also the rotational dark matter flows may influence the angular momentum transport and thus the evolving cosmic structures in a non-trivial way. In the words of Adamek et al. (2016), this indicates how spacetime is dragged around by vortical matter flows.
@@ 5.4 Time evolution
Compared to the right panel of Figure 5 in Barrera-Hinojosa et al. (2021), where the power spectrum of the vector potential is found to decrease over time, from $z=1.5$ to $z=0$, the reader should keep in mind that there is a factor $a^{2}$ difference between the adopted conventions for the power spectrum of $\mathbf{B}$. Moreover, in Barrera-Hinojosa et al. (2021), the ratio is found to increase towards the smallest scales, in contrast with our results. This is most likely due to the Adaptive Mesh Refinement technique implemented in GRAMSES, which enables a much higher spatial resolution than DTFE. It is still unclear, however, whether this increase manifests as a result of spurious power introduced by the AMR or rather as a pure GR indication.
Since the non-linear prediction of the vector potential power spectrum $\Delta_{\mathbf{B}}^{\mathrm{n.l.}}$ gives overall a good estimate for the fully non-linear vector potential, and its evolution, relevant information can already be retrieved from this quantity. In order to highlight the redshift dependence of the power spectrum independently of its overall amplitude at a given redshift due to the linear growth of structure, we normalise the power spectrum $\Delta_{\mathbf{B}}^{\mathrm{n.l.}}$ by the analytic growth factor of $\Delta_{\mathbf{B}}^{\mathrm{PT}}$, which is (see Equation B.4 of Adamek et al., 2014)
@@
is the growth factor for the scalar potential and $g_{\infty}$ is chosen so that $g(0)=1$ (see Equation A.4 of Lu et al., 2009). The left panel of Figure 18 shows $\Delta_{\mathbf{B}}^{\mathrm{n.l.}}/f_{\mathbf{B}}^{\mathrm{PT}}$ for different redshifts.
@@
Here, we can clearly see how the non-linear growth of vector modes is highly scale-dependent and how the power is transferred over time. As expected, the non-linearity in the power spectrum of the gravito-magnetic vector potential is observed to increase with redshift from $z=20$ to $z=0$. From the right panel, one can quantify the increase of the amplitude of the non-linear approximation with respect to perturbation theory. In particular, at $z=0$, the non-linear amplitude is $\approx 175$ times larger than PT, for scales close to k=100hk=100\;h Mpc-1, while being $\approx 75$ times larger for k=10hk=10\;h Mpc-1.
@@
For the smallest scales, in the range $10\text{--}100\;h$ Mpc-1, the pure non-linear evolution is intensively boosted from $z=1\text{--}2$ onwards. However, as shown in the right panel and already stated at the beginning of this section, the ratio between the non-linear approximation for the power spectrum of the vector potential and the non-linear analogue for the scalar potential remains of the order of $10^{-5}$. Contrary to Thomas et al. (2015b), where this ratio is not monotonic over the redshift range (see Fig. 9 of the reference), here instead we find a monotonic increase over time. In particular, for the smallest scales, this ratio starts growing rapidly since $z=1\text{--}2$. Therefore, we can conclude that there is no indication pushing towards the direction of a relevant gravito-magnetic effect arising during a particular phase in the evolution of structures, other than the rapidly accelerated yet continuous growth of non-linearity at low redshift.
@@ 6 Conclusions
On non-linear scales, cosmic structure formation is typically studied through Newtonian $N$-body simulations, while relativistic effects are only considered at large scales, of the order of the Hubble horizon, using perturbative methods, e.g. to study primordial non-gaussianity signatures in galaxy surveys. The post-Friedmann approach provides a unified framework for small and large scales that generalises the weak-field approximation to cosmology. In this framework, at the leading order in the $c^{-n}$ expansion, namely in the Newtonian limit of small peculiar velocities, a frame-dragging effect is expected to arise in terms of a gravito-magnetic vector potential. This potential is sourced by the transverse part of the momentum field and is thus expected to be non-zero if measured from Newtonian $N$-body simulations, as confirmed by previous works on cosmological scales (Bruni et al., 2014; Thomas et al., 2015b).
The IllustrisTNG project provides a suite of simulations that offer a unique opportunity to investigate this gravito-magnetic effect not only on cosmological scales but also in the highly non-linear regime, down to galactic scales, where structures have fully collapsed, have gone through the virialisation process and the vorticity that has been generated in the multi-stream regime becomes significant (Pueblas and Scoccimarro, 2009). Our goal here has been the investigation of gravito-magnetic effects on these previosly unexplored smaller non-linear scales, using the post-Friedmann framework on simulations of the IllustrisTNG project.
To this aim, three dark-matter-only simulations have been considered, i.e. TNG300-2-Dark, TNG100-2-Dark, and TNG50-2-Dark, which sample different non-linear regimes in the formation of structures. Following the methodology developed by Bruni et al. (2014) and Thomas et al. (2015a), the source fields of the gravito-magnetic potential have been extracted via the Delaunay Tessellation Field Estimator, that provides a detailed reconstruction of the density and velocity fields, from which power spectra were computed.
@@
In our results the power spectra of the velocity gradients show the onset of non-linearity and the generation of vorticity. At around $k\approx 1.5\;h$ Mpc-1 for the $205$ and $75$ $h^{-1}$ Mpc simulations, and at around $k\approx 3\;h$ Mpc-1 for the $35$ $h^{-1}$ Mpc simulation, the velocity divergence shows a dip while the vorticity exhibits a peak and start to dominate at smaller scales as the system transitions into the non-linear regime, in agreement with previous studies (Pueblas and Scoccimarro, 2009; Jelic-Cizmek et al., 2018; Barrera-Hinojosa et al., 2021). This reveals that rotational (i.e. divergence-free) modes emerge naturally in cosmological simulations as a consequence of orbit crossing in collapsing regions, a process that becomes increasingly important as structures evolve beyond the single-stream approximation due to the collision-less nature of dark matter.
@@
The simulation results are compared with both second-order perturbation theory and a non-linear approximation. This work shows for the first time that the non-linear prediction provides actually a good estimate for the non-linear power spectrum of the vector potential. A correction for the missing power in the gravito-magnetic vector potential power spectrum, inspired by a technique used in the Sunyaev-Zel’dovich effect literature (Park et al., 2018) and computed using the non-linear matter power spectrum from HaloFit, confirms that at large scales, the corrected spectra match perturbation theory predictions, while at small scales, the non-linear approximation remains a valid estimate. This also confirms that the correction employed is actually accurate in reproducing consistent results across simulations of different size.
The ratio of the vector to scalar potential, corrected for the missing power, is found to be $\approx 2\text{--}4\times 10^{-5}$ across all simulations, consistent with previous studies (Bruni et al., 2014; Thomas et al., 2015b; Adamek et al., 2016; Barrera-Hinojosa et al., 2021) and second-order perturbation theory (Lu et al., 2009). Interestingly, the non-linear prediction for the vector power spectrum is found to reach an approximately constant value of $\approx 5\times 10^{-5}$ at very small scales, which is only slightly larger than expected in 2nd-order PT, i.e. $\approx 3\times 10^{-5}$.
Furthermore, differently from Bruni et al. (2014) and Thomas et al. (2015b), the global solution for the gravito-magnetic potential over the entire simulation volume is also reconstructed from spectral analysis. Here, a visual representation of its magnitude is presented for the first time for Newtonian $N$-body simulations. The gravito-magnetic potential, while weaker, is generally correlated with the scalar potential. It also shows clear non-local structures and organised vortical patterns that indicate a coupling between rotational flows and the dragging of spacetime.
@@
However, the results obtained here for the gravito-magnetic potential have also been confirmed by cosmological studies performed with the fully relativistic code GRAMSES (Barrera-Hinojosa et al., 2021). Contrary to the case of the post-Friedmann approach, which is restricted to weak-field regime approximations where relativistic terms are treated as small perturbations to the background FLRW metric, GRAMSES adopts the 3+1 formalism and the scalar and vector modes of the gravitational field are treated fully non-linearly. The code makes also use of the Adaptive Mesh Refinement in high-density regions so that it can reach spatial resolutions down to $2\;h^{-1}$ kpc. Here, indeed, a simulation of $256\;h^{-1}$ Mpc with mass resolution of $1.33\times 10^{9}\;h^{-1}\,\mathrm{M}_{\odot}$ was found to give a ratio between the gravito-magnetic and scalar potentials of the order of $10^{-3}$ inside dark matter halos of $10^{12.5}\text{--}10^{15}\;h^{-1}\,\mathrm{M}_{\odot}$. However, while this contribution remains small on these scales, the actual computation of the frame-dragging for simulations of the size of TNG50-2-Dark, having a mass resolution of $2.9\times 10^{6}\;h^{-1}\,\mathrm{M}_{\odot}$, is yet to be performed using GRAMSES.
Importantly, as current and previous studies have focused exclusively on dark-matter-only simulations, the impact of baryons on such small scales still remains to be assessed. The presence of baryons will affect the source term for the gravito-magnetic field, which can be broken down into three components as in Equation (28), from left to right: vorticity, density-weighted vorticity and $\mathbf{\nabla}\delta\times\mathbf{v}$. It has been shown that when baryonic physics is included in simulations, there is a suppression of the matter clustering on small scales due to processes such as AGN feedback (vanDaalen:2011; Saha:2024). On the flip side, baryons contribute to the generation of vorticity via feedback-driven flows, shocks, and pressure gradients. Hence, the net effect on the rotational component of the momentum field of matter, and hence on the gravito-magnetic potential, is not a priori clear. However, since Figure 4 shows that on small scales the $\mathbf{\nabla}\delta\times\mathbf{v}$ term is larger than the density-weighted vorticity term, it is likely that the net effect is a suppression of the gravito-magnetic field.
@@
WB and MC are indebted to the Italian Space Agency (ASI) for their continuing support through contract 2018-24-HH.0 and its addendum 201824-HH.1-2022 to the National Institute for Astrophysics (INAF). CB-H is supported by the Chilean National Agency of Research and Development (ANID) through grant FONDECYT/Postdoctorado No. 3230512. MB has been supported by UK STFC Grants No. ST/S000550/1.